OpenStudy (anonymous):

what are the solutions for x? 3x^2=-9x

OpenStudy (p0sitr0n):

x = -3 x = 0 3x2 + 9x = 0 use quadratic formula to calculate x1 and x2

OpenStudy (p0sitr0n):

\[x1,x2 = (-b \pm \sqrt{b^2-4ac})/2a\] a = 3 b=9 c = 0 if 2 roots = two intersection points for the parabola 1 root = one 0 = no roots, parabola doesnt intercept the axis